If a bicyclist can drive down a one minute hill three times as fast as she can ride back up the hill, the round trip takes 20 minutes, how fast does she drive down the hill in miles per hour.
Answer by richwmiller(17219)

What is a one minute hill?
Do you mean one mile hill?
r*t=d
3x*t=1
x*(20-t)=1
t=5 x=1/15
3/15*5=1
1/15(15)=1
5 minutes going down and 15 going up
1/5 mph going down and 1/15 going up the hill.